1. The elderly glued their eyes on a TV 6 hours/day. The TV is sourced from a 220 Volts and the I flows through the TV is 0.5 Amp. If the electric company charges Php 0.092 per kWh, then the cost of using electric energy for TV for 30 days.

a) What is the power of TV? P = V I

b) What is the electric energy of TV? electric power x time interval

c) What is the cost of using electric energy for 30 days? electric energy x cost

Given: Time interval = 6 hours x 30 = 180 hours

Voltage (V) = 220 Volt

Electric current (I) = 0.5 Ampere


1
Expert's answer
2021-03-08T08:16:14-0500

(a)

"P=VI=220\ V\cdot0.5\ A=110\ W."

(b)

"E=Pt=110\ W\cdot6\ \dfrac{h}{day}\cdot30\ days=19.8\ kWh."

(c)

"Cost=19.8\ kWh\cdot 0.092\ \dfrac{Php}{kWh}=Php\ 1.82."